8) A 4.00-kg mass hangs vertically from a massless spring. The spring constant of the spring is 100 N/m. The mass is pulled d do not solve using shm
Answer #1

please upvote ? if you like my answer ?

When mass is pulled , a za max 0.02m Thes extension causes development of elastic potential Evergy When mass is 0.01 m. abov

energy remained conseurd so I kurn - Ik x 2 + 1 m ² v² t error x²) = 100 (0.02² -0.01%) = 0.075 m²/2 = v=0.0866 m/s option A